OpenStudy (anonymous):

A text font fits 12 characters per inch. Using the same font, how many characters can be expected per yard of text?

OpenStudy (anonymous):

how many inches in a foot?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

how many feet in a yard?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

what is \(12\times 12\times 3\) ?

OpenStudy (anonymous):

so then mulitiply 12x3

OpenStudy (kobeni-chan):

there are 3 ft per yard and 12 inchs per foot, 36 characters

OpenStudy (anonymous):

ohh oops lol

OpenStudy (anonymous):

no 36 inches in a yard, not 36 characters multiply \(36\times 12\)